OpenStudy (anonymous):

SORRY>>> RIGHT QUESTIONS Which of the following is a pair of equivalent fractions? 8/12 and 16/48 3/14 and 6/42 4/24 and 8/48 3/8 and 11/16

OpenStudy (barrycarter):

Hint: reduce all fractions to lowest terms

OpenStudy (anonymous):

Its 4/24 and 8/48 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{4}{24}=\frac{4}{24}*1=\frac{4}{24}*\frac{2}{2}=\frac{8}{48}\]